A right-angle triangle has hypoteneuse 10 inches, and the height drawn from the right angle to the hypoteneuse is 6 inches. Find the area of this right-angle triangle
Answer by usyim88hk(158)

First we should find the third side.
Let x be the length of the third side
10^2 = 6^2 + x^2
100 = 36 + x^2
64 = x^2
x = 8
So the third side is eight.
And now we can use the two sides other then the hypoteneuse to find the area of this triangle
8*6/2
= 24